OpenStudy (anonymous):

Find the value of the sec 19° using your calculator.

OpenStudy (anonymous):

-.- funny

OpenStudy (anonymous):

So you thought we were just a calculator?

OpenStudy (anonymous):

No I just don't understand the question. No need to take this personally @wio

OpenStudy (anonymous):

Okay, well do you know that \[\cos(19^\circ)=\frac{1}{\sec(19^\circ)}\]So that means \[\sec(19^\circ)=\frac{1}{\cos(19^\circ)}\]Can you do it now?

OpenStudy (anonymous):

Okay so I get it now. So it would be 1.05, right?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

Okay! Thank you for the help!!

OpenStudy (anonymous):

@wio can you help me with another I don't understand?

OpenStudy (anonymous):

Just post a new question. Makes it easier to keep track of stuff.

OpenStudy (anonymous):

okay